OpenStudy (anonymous):

find the measure of an angle whose supplement measures seventeen times its measure

zepdrix (zepdrix):

Hmm so let's see. If angles are supplementary, they add up to \(\large 180^o\) right? `find the measure of an angle`, let's call this angle \(\large x\). `whose supplement`, hmm let's come up with an equation for the supplement. I guess we can call it.... \(\large s\) maybe..? :P If x and s are supplementary, they add up to 180. \(\large x+s=180^o\) So if we want an equation for the supplementary angle, we can write it as, \(\large 180^o-x=s\) `whose supplement measures seventeen times its measure` this wording is really confusing. I guess what they're saying is, who's supplement equals 17 times the original angle measure. Which we can write like this \(\large s=17x\). And we're trying to solve for x.

OpenStudy (anonymous):

ok. I see let's see if I can figure it out now.

OpenStudy (anonymous):

so x = 18 degrees

zepdrix (zepdrix):

\[\large \color{green}{s=180^o-x}\] \[\large \color{green}{s}=17x \qquad \rightarrow \qquad \color{green}{180-x}=17x\] Which simplifies down to,\[\large 180=18x \qquad \rightarrow \qquad x=\frac{180}{18}\] Do these steps look the same as yours? I think maybe you just messed up the division at the end :o

OpenStudy (anonymous):

Yup I messed up division its 10. Thanks. This problem is confusing though. I will try some similar ones and see if I can do them based on this example. Thanks again!

zepdrix (zepdrix):

Cool c: